OpenStudy (anonymous):

Solve the equation. 1/4 - t/2 = 15/8

OpenStudy (anonymous):

\[2-4t=15\] is a good start

OpenStudy (anonymous):

t=-13/4

OpenStudy (anonymous):

t/2 = 1/4 - 15/8 = (2 - 15)/8 t = -13/4